# Imitation Learning from a Single Temporally Misaligned Video

### v1

@@ Abstract
We examine the problem of learning sequential tasks from a single visual demonstration. A key challenge arises when demonstrations are temporally misaligned due to variations in timing, differences in embodiment, or inconsistencies in execution. Existing approaches treat imitation as a distribution-matching problem, aligning individual frames between the agent and the demonstration. However, we show that such frame-level matching fails to enforce temporal ordering or ensure consistent progress. Our key insight is that matching should instead be defined at the level of sequences. We propose that perfect matching occurs when one sequence successfully covers all the subgoals in the same order as the other sequence. We present ORCA (ORdered Coverage Alignment), a dense per-timestep reward function that measures the probability of the agent covering demonstration frames in the correct order. On temporally misaligned demonstrations, we show that agents trained with the ORCA reward achieve $4.5$x improvement ($0.11\rightarrow 0.50$ average normalized returns) for Meta-world tasks and $6.6$x improvement ($6.55\rightarrow 43.3$ average returns) for Humanoid-v4 tasks compared to the best frame-level matching algorithms. We also provide empirical analysis showing that ORCA is robust to varying levels of temporal misalignment. Our code is available at https://github.com/portal-cornell/orca/
@@ 1 Introduction
Inverse reinforcement learning (IRL) (Abbeel & Ng, 2004; Ziebart et al., 2008) provides a principled way to infer rewards by matching the learner’s trajectory to expert demonstrations. Recent IRL approaches apply optimal transport (OT) (Peyré & Cuturi, 2020) to align visual embeddings between frames (Cohen et al., 2022; Haldar et al., 2023b, a; Guzey et al., 2024; Tian et al., 2024; Liu et al., 2024; Fu et al., 2024b). However, these methods operate at the frame level, ignoring temporal dependencies: an agent can revisit subgoals, skip ahead, or stall without penalty. As a result, frame-matching approaches fail on sequence-matching tasks, where the correct subgoal order is crucial for success.
@@ 3 Desiderata of Sequence-Matching Rewards
IRL can be formulated as a distribution matching problem between the learner and the demonstrator on “moments”, i.e., expectations of reward basis functions (Swamy et al., 2021). Hence, solving IRL is equivalent to optimizing Integral Probability Metrics (IPMs) between the learner and demonstrator distributions (Sun et al., 2019). Recent works use optimal transport (OT) by minimizing the Wasserstein distance (an IPM) between embeddings of the frames in the learner and demonstration trajectories. They work with Markovian moments that depend on only a single frame.
@@ 3.1 Optimal Transport Fails to Respect Ordering
Prior works in imitation learning use Optimal Transport (OT) (Peyré & Cuturi, 2020), specifically the Wasserstein distance, to measure how closely a learner trajectory matches a demonstration trajectory (Papagiannis & Li, 2022; Tian et al., 2024; Fu et al., 2024b; Kedia et al., 2024). These approaches assume a Markovian distance metric $d(\cdot,\cdot)$ that measures the distance between the embeddings of a learner and demonstration frame. Given a learner trajectory $\xi=\{o_{t}\}^{T}_{t=1}$ and a video demonstration $\tilde{\xi}=\{\tilde{o}_{j}\}^{\tilde{T}}_{j=1}$, they define corresponding learner and demonstration distributions that are uniform in time. Specifically, $\rho=\frac{1}{T}\sum^{T}_{t=1}\delta_{o_{t}}$, where $\delta_{o_{t}}$ is a Dirac distribution centered on $o_{t}$. The demonstration similarly has a distribution of $\tilde{\rho}=\frac{1}{\tilde{T}}\sum^{\tilde{T}}_{j=1}\delta_{\tilde{o}_{j}}$. Consequently, the matching matrices are evenly weighted, defined as the set $M=\{\mu\in\mathbb{R}^{T\times\tilde{T}}:\mu\mathbf{1}=\frac{1}{T},\mu^{T}\mathbf{1}=\frac{1}{\tilde{T}}\}$. With an entropy regularizer $\mathcal{H(\cdot)}$, OT solves for the optimal matching matrix:
@@ 3.3 TemporalOT Fails to Cover All Subgoals under Temporal Misalignment
Recognizing the limitations of OT, Fu et al. (2024b) propose TemporalOT, which alters the OT objective to optimize over a temporally masked cost matrix. Specifically, the mask is a variant of the diagonal matrix with a hyperparameter $k_{w}$ to define the width of the diagonal:
@@ 4.1 ORCA: Ordered Coverage Reward Function
  $P_{i,j}=\exp(-\lambda d(o_{t},\tilde{o}_{j}))\text{ for }t\in[[T]],j\in[[\tilde{T}]]$
@@
  Return Rewards $R_{\text{ORCA}}=\{C_{t,\tilde{T}-1}P_{t,\tilde{T}}\mid t\in[[T]]\}$
@@ 4.2 Analysis
Let $\xi^{-}$ be a trajectory that is out of order; specifically, there exists a subgoal $\tilde{o}_{j}$ such that $\tilde{o}_{j}$ is occupied at time $t$ but $\tilde{o}_{j-1}$ is not yet occupied. Let $\xi^{+}$ be a trajectory that is identical to $\xi^{-}$, except that it occupies $\tilde{o}_{j-1}$ before time $t$. Then, $\mathcal{R}_{\text{ORCA}}(o^{+}_{t},\tilde{\xi})>\mathcal{R}_{\text{ORCA}}(o^{-}_{t},\tilde{\xi})$.
@@
Let $\xi^{-}$ be a trajectory that occupies $\tilde{o}_{j-1}$ at time $t-1$ and continues to occupy $\tilde{o}_{j-1}$ at time $t$, instead of progressing towards $\tilde{o}_{j}$. Let $\xi^{+}$ be an identical trajectory that progresses towards $\tilde{o}_{j}$ at $t$, and assume that neither trajectory has been closer to $\tilde{o}_{j}$ before. $\mathcal{R}_{ORCA}(o^{+}_{t},\tilde{\xi})>\mathcal{R}_{ORCA}(o^{-}_{t},\tilde{\xi})$.
@@ 4.3 Pretraining
Table 2: Humanoid results on temporally misaligned demonstrations. Results are presented as the mean returns with standard error. TemporalOT is abbreviated to TOT. For results of all baselines, see Table 4 in Appendix C.4.
@@ 5.1 Experimental Setup
  • Meta-World (Yu et al., 2021). Following Fu et al. (2024b), we use ten tasks from the Meta-world environment to evaluate the effectiveness of ORCA reward in the robotic manipulation domain. We classify the tasks into three difficulty levels based on the number of types of motions required and whether the task requires precision. Visual demonstrations are generated using hand-engineered policies provided by the environment.
@@
RL Policy. For Meta-world, we follow the RL setup in  Fu et al. (2024b). We train DrQ-v2 (Yarats et al., 2021) with state-based input for 1M steps and evaluate the policy every 10k steps on 10 randomly seeded environments. For the Humanoid environment, we train SAC (Haarnoja et al., 2018) for 2M steps and evaluate the policy every 20k steps on 8 environments. All policies use state-based input, and in metaworld we include an additional feature that represents the percentage of total timesteps passed. Appendix C.1 contains RL training details and hyperparameters.
Baselines. We compare ORCA against baselines that use frame-level matching algorithms: OT (Tian et al., 2024), TemporalOT (Fu et al., 2024b), and DTW (Sakoe & Chiba, 1978). We also compare ORCA, which is first trained on TemporalOT rewards for half of the total timesteps, then ORCA rewards for the remaining timesteps, against ORCA(NP), which fully trains on ORCA rewards without any initialization.ompared to All approaches use the pretrained ResNet50 (He et al., 2015) to extract visual features and cosine similarity as the distance function. We also include a simple baseline Threshold, which tracks the subgoals completed based on a threshold distance, and a transformer-based approach RoboCLIP (Sontakke et al., 2024), which directly encodes an entire video. Details in Appendix C.2.
@@ 5.2 How well does ORCA perform given temporally misaligned demonstrations?
Metaworld. For each original demonstration, we subsample it such that the first one-fifth retains the original execution speed, but the rest is sped up by five to ten times. In Table 1, the ORCA reward significantly outperforms other baselines and trains agents to acquire the highest average normalized return of $0.50$. The Push task is difficult for all approaches and the only one where ORCA or ORCA(NP) does not achieve top performance. As analyzed by Fu et al. (2024c), the distance matrix is generally noisier for the Push task due to the small size of the target object, which benefits a sparser reward function like Threshold. Meanwhile, rewards based on frame-level matching algorithms perform poorly on all tasks, revealing their weakness when encountering temporal misalignment.
@@
Humanoid. Because there does not exist an expert policy for the Humanoid tasks, we generate demonstrations by interpolating ten frames between the initial and final joint positions. These demonstrations are naturally temporally misaligned because the environment is unstable, making it impossible to follow the subgoals at the same speed. Table 2 shows that agents trained with ORCA(NP) achieve the highest average cumulative return of $62.9$. Due to poor TemporalOT performance, ORCA does not benefit from pretraining, and ORCA(NP) thus performs better because it is trained on the ordered coverage reward for more steps. We further investigate the failure of TemporalOT in Sec. 5.4. Meanwhile, Fig. 14 in Appendix C.4 shows how the ORCA reward satisfies Prop. 4.1 and 4.2, covering the subgoals as quickly as possible and successfully completing the task.
@@
Varying Misalignment Level. We identify two types of temporal misalignment: either a slower demonstration that contains pauses, or a faster demonstration that accelerates through a segment. For each misalignment type, we randomly perturb the original demonstrations of three Meta-world tasks (Door-open, Window-open, Lever-pull), where the misalignment level controls how varied and nonlinear speed changes are. See appendix C.3 for details. In Fig. 4, ORCA consistently maintains a higher return compared to TemporalOT as the demonstrations become more misaligned. Meanwhile, TemporalOT’s performance significantly deteriorates when there is any level of misalignment. When the demonstrations are sped up, because TemporalOT encourages agents to spend an equal amount of time at each subgoal (in-depth discussion in Sec 5.4), TemporalOT agents often cannot finish the task in time. This problem is further exacerbated when the demonstrations are slowed down and longer than the learner trajectory. ORCA’s performance also worsens more given slower demonstrations compared to faster ones. In addition to being affected by poor initialization due to TemporalOT’s poor performance, we observe that when ORCA agents fail, they successfully follow the general motions of the demonstration, but they miss details (e.g., aligning the gripper with the target object). We hypothesize that this behavior is caused by the frame-level distance metric, which pays more attention to the general robot arm motions than details, allowing most subgoals to achieve relatively good coverage.
5.3 How does failure to enforce subgoal ordering affect performance?
@@
Figure 4: Results given varying levels of temporal misalignment. We report the mean expert-normalized returns with standard error across 3 Meta-world tasks. We generate 3 perturbed demonstrations per task per misalignment level (L=Low, H=High).
Figure 5: Average expert-normalized returns across all Meta-world tasks on temporally aligned demonstrations.
5.4 How does failure to enforce full subgoal coverage affect performance?
@@ 5.5 How does pretraining affect ORCA’s performance?
Pretraining leads to better ORCA performance when the pretraining strategy is able to obtain some success. In Table. 1, ORCA is equal or better than ORCA(NP) on temporally misaligned demonstrations. In Humanoid tasks, ORCA(NP) achieves higher overall performance because TemporalOT almost entirely fails on every task. The effect of pretraining is thus most apparent when the demonstrations are temporally aligned, where TemporalOT performs well because the setting satisfies its assumption. In Fig. 5, ORCA achieves an average normalized return of $0.57$ compared to ORCA(NP) ($0.33$) on aligned demonstrations.

### v2

@@ Abstract
We examine the problem of learning sequential tasks from a single visual demonstration. A key challenge arises when demonstrations are temporally misaligned due to variations in timing, differences in embodiment, or inconsistencies in execution. Existing approaches treat imitation as a distribution-matching problem, aligning individual frames between the agent and the demonstration. However, we show that such frame-level matching fails to enforce temporal ordering or ensure consistent progress. Our key insight is that matching should instead be defined at the level of sequences. We propose that perfect matching occurs when one sequence successfully covers all the subgoals in the same order as the other sequence. We present ORCA (ORdered Coverage Alignment), a dense per-timestep reward function that measures the probability of the agent covering demonstration frames in the correct order. On temporally misaligned demonstrations, we show that agents trained with the ORCA reward achieve $4.5$x improvement ($0.11\rightarrow 0.50$ average normalized returns) for Meta-world tasks and $6.6$x improvement ($6.55\rightarrow 43.3$ average returns) for Humanoid-v4 tasks compared to the best frame-level matching algorithms. We also provide empirical analysis showing that ORCA is robust to varying levels of temporal misalignment. The project website is at https://portal-cornell.github.io/orca/
@@ 1 Introduction
Inverse reinforcement learning (IRL) (Abbeel & Ng, 2004; Ziebart et al., 2008) provides a principled way to infer rewards by matching the learner’s trajectory to expert demonstrations. Recent IRL approaches apply optimal transport (OT) (Peyré & Cuturi, 2020) to align visual embeddings between frames (Cohen et al., 2022; Haldar et al., 2023b, a; Guzey et al., 2024; Tian et al., 2024; Liu et al., 2024; Fu et al., 2024c). However, these methods operate at the frame level, ignoring temporal dependencies: an agent can revisit subgoals, skip ahead, or stall without penalty. As a result, frame-matching approaches fail on sequence-matching tasks, where the correct subgoal order is crucial for success.
@@ 3 Desiderata of Sequence-Matching Rewards
IRL can be formulated as a distribution matching problem between the learner and the demonstrator on “moments", i.e., expectations of reward basis functions (Swamy et al., 2021). Hence, solving IRL is equivalent to optimizing Integral Probability Metrics (IPMs) between the learner and demonstrator distributions (Sun et al., 2019). Recent works use optimal transport (OT) by minimizing the Wasserstein distance (an IPM) between embeddings of the frames in the learner and demonstration trajectories. They work with Markovian moments that depend on only a single frame.
@@ 3.1 Optimal Transport Fails to Respect Ordering
Prior works in imitation learning use Optimal Transport (OT) (Peyré & Cuturi, 2020), specifically the Wasserstein distance, to measure how closely a learner trajectory matches a demonstration trajectory (Papagiannis & Li, 2022; Tian et al., 2024; Fu et al., 2024c; Kedia et al., 2025). These approaches assume a Markovian distance metric $d(\cdot,\cdot)$ that measures the distance between the embeddings of a learner and demonstration frame. Given a learner trajectory $\xi=\{o_{t}\}^{T}_{t=1}$ and a video demonstration $\tilde{\xi}=\{\tilde{o}_{j}\}^{\tilde{T}}_{j=1}$, they define corresponding learner and demonstration distributions that are uniform in time. Specifically, $\rho=\frac{1}{T}\sum^{T}_{t=1}\delta_{o_{t}}$, where $\delta_{o_{t}}$ is a Dirac distribution centered on $o_{t}$. The demonstration similarly has a distribution of $\tilde{\rho}=\frac{1}{\tilde{T}}\sum^{\tilde{T}}_{j=1}\delta_{\tilde{o}_{j}}$. Consequently, the matching matrices are evenly weighted, defined as the set $M=\{\mu\in\mathbb{R}^{T\times\tilde{T}}:\mu\mathbf{1}=\frac{1}{T},\mu^{T}\mathbf{1}=\frac{1}{\tilde{T}}\}$. With an entropy regularizer $\mathcal{H(\cdot)}$, OT solves for the optimal matching matrix:
@@ 3.3 TemporalOT Fails to Cover All Subgoals under Temporal Misalignment
Recognizing the limitations of OT, Fu et al. (2024c) propose TemporalOT, which alters the OT objective to optimize over a temporally masked cost matrix. Specifically, the mask is a variant of the diagonal matrix with a hyperparameter $k_{w}$ to define the width of the diagonal:
@@ 4.1 ORCA: Ordered Coverage Reward Function
Runtime complexity. ORCA’s runtime complexity is $\mathcal{O}(T\cdot\tilde{T})$. This is the lower bound time complexity of any method that relies on a frame-level distance matrix, including all OT-based methods. To validate this, we experimentally tested the latency of ORCA against a variety of common baselines, finding that ORCA is faster than TemporalOT and comparable to OT. See Appendix C.6 for details.
@@
  $P_{t,j}=\exp(-\lambda d(o_{t},\tilde{o}_{j}))\text{ for }t\in[[T]],j\in[[\tilde{T}]]$
@@
  Return Rewards $\mathcal{R}_{\text{ORCA}}=\{C_{t,\tilde{T}-1}P_{t,\tilde{T}}\mid t\in[[T]]\}$
@@ 4.2 Analysis
Let $\xi^{-}$ be a trajectory that is out of order; specifically, there exists a subgoal $\tilde{o}_{j}$ such that $\tilde{o}_{j}$ is occupied at time $t$ but $\tilde{o}_{j-1}$ is not yet occupied. Let $\xi^{+}$ be a trajectory that is identical to $\xi^{-}$, except that it occupies $\tilde{o}_{j-1}$ before time $t$. Then, $\mathcal{R}_{\text{ORCA}}(o^{+}_{t},\xi^{+},\tilde{\xi})>\mathcal{R}_{\text{ORCA}}(o^{-}_{t},\xi^{-},\tilde{\xi})$.
@@
Let $\xi^{-}$ be a trajectory that occupies $\tilde{o}_{j-1}$ at time $t-1$ and continues to occupy $\tilde{o}_{j-1}$ at time $t$, instead of progressing towards $\tilde{o}_{j}$. Let $\xi^{+}$ be an identical trajectory that progresses towards $\tilde{o}_{j}$ at $t$, and assume that neither trajectory has been closer to $\tilde{o}_{j}$ before. Then, $\mathcal{R}_{ORCA}(o^{+}_{t},\xi^{+},\tilde{\xi})>\mathcal{R}_{ORCA}(o^{-}_{t},\xi^{-},\tilde{\xi})$.
@@ 4.3 Pretraining
Table 2: Humanoid results on temporally misaligned demonstrations. Results are presented as the mean returns with standard error. TemporalOT is abbreviated to TOT. For results of all baselines, see Table 4 in Appendix C.
@@ 5.1 Experimental Setup
  • Meta-World (Yu et al., 2020). Following Fu et al. (2024c), we use ten tasks from the Meta-world environment to evaluate the effectiveness of ORCA reward in the robotic manipulation domain. On each rollout, the locations of task-relevant objects are randomized according to the default setup of Metaworld. The training and evaluation seeds are also different. We classify the tasks into three difficulty levels based on the number of types of motions required and whether the task requires precision. Visual demonstrations are generated using hand-engineered policies provided by the environment.
@@
RL Policy. For Meta-world, we follow the RL setup in  Fu et al. (2024c). We train DrQ-v2 (Yarats et al., 2021) with state-based input for 1M steps and evaluate the policy every 10k steps on 10 randomly seeded environments. For the Humanoid environment, we train SAC (Haarnoja et al., 2018) for 2M steps and evaluate the policy every 20k steps on 8 environments. All policies use state-based input, and in metaworld we include an additional feature that represents the percentage of total timesteps passed. Appendix B.4 contains RL training details and hyperparameters.
Baselines. We compare ORCA against baselines that use frame-level matching algorithms: OT (Tian et al., 2024), TemporalOT (Fu et al., 2024c), and DTW (Sakoe & Chiba, 1978). We also compare ORCA, which is pretrained on TemporalOT rewards for half of the total timesteps and ORCA rewards for the remaining timesteps, against ORCA(NP), which fully trains on ORCA rewards without any initialization. All approaches use the pretrained ResNet50 (He et al., 2016) to extract visual features and cosine similarity as the distance function. We include a simple baseline Threshold, which tracks the subgoals completed based on a threshold distance, and a transformer-based approach RoboCLIP (Sontakke et al., 2024), which directly encodes an entire video. Details are in Appendix B.5. We also investigate a language-conditioned baseline and a traditional IRL baseline on a subset of the tasks in Appendix C.3.
@@ 5.2 How well does ORCA perform given temporally misaligned demonstrations?
Metaworld. For each original demonstration, we subsample it such that the first one-fifth retains the original execution speed, but the rest is sped up by five to ten times. In Table 1, the ORCA reward significantly outperforms other baselines and trains agents to acquire the highest average normalized return of $0.50$. The Push task is difficult for all approaches and the only one where ORCA or ORCA(NP) does not achieve top performance. As analyzed by Fu et al. (2024d), the distance matrix is generally noisier for the Push task due to the small size of the target object, which benefits a sparser reward function like Threshold. Meanwhile, rewards based on frame-level matching algorithms perform poorly on all tasks, revealing their weakness when encountering temporal misalignment.
@@
Humanoid. Because there does not exist an expert policy for the Humanoid tasks, we generate demonstrations by interpolating ten frames between the initial and final joint positions. These demonstrations are naturally temporally misaligned because the environment is unstable, making it impossible to follow the subgoals at the same speed. Table 2 shows that agents trained with ORCA(NP) achieve the highest average cumulative return of $62.9$. Due to poor TemporalOT performance, ORCA does not benefit from pretraining, and ORCA(NP) thus performs better because it is trained on the ordered coverage reward for more steps. We further investigate the failure of TemporalOT in Sec. 5.4. Meanwhile, Fig. 15 in Appendix C shows how the ORCA reward satisfies Prop. 4.1 and 4.2, covering the subgoals as quickly as possible and successfully completing the task.
Varying Misalignment Level. We identify two types of temporal misalignment: either a slower demonstration that contains pauses, or a faster demonstration that accelerates through a segment. For each misalignment type, we randomly perturb the original demonstrations of three Meta-world tasks (Door-open, Window-open, Lever-pull), where the misalignment level controls how varied and nonlinear speed changes are. See appendix B.6 for details.
@@
In Fig. 4, ORCA consistently maintains a higher return compared to TemporalOT as the demonstrations become more misaligned. Meanwhile, TemporalOT’s performance significantly deteriorates when there is any level of misalignment. When the demonstrations are sped up, because TemporalOT encourages agents to spend an equal amount of time at each subgoal (in-depth discussion in Sec 5.4), TemporalOT agents often cannot finish the task in time. This problem is further exacerbated when the demonstrations are slowed down and longer than the learner trajectory. ORCA’s performance also worsens more given slower demonstrations compared to faster ones.
In addition to being affected by poor initialization due to TemporalOT’s poor performance, we observe that when ORCA agents fail, they successfully follow the general motions of the demonstration, but they miss details (e.g., aligning the gripper with the target object). We hypothesize that this behavior is caused by the frame-level distance metric, which pays more attention to the general robot arm motions than details, allowing most subgoals to achieve relatively good coverage.
5.3 How important is enforcing subgoal ordering?
@@
Figure 4: Results given varying levels of temporal misalignment. We report the mean expert-normalized returns with standard error across 3 Meta-world tasks (Door open, Window open, Lever pull). We generate 3 perturbed demonstrations per task per misalignment level (L=Low, H=High), training on each demonstration separately.
Figure 5: Results given temporally aligned demonstrations. We report mean expert-normalized returns with standard error across all 10 Meta-world tasks.
5.4 How important is enforcing full subgoal coverage?
@@ 5.5 How does pretraining affect ORCA’s performance?
Pretraining leads to better ORCA performance when the pretraining strategy is able to obtain some success. In Table. 1, ORCA is equal or better than ORCA(NP) on temporally misaligned demonstrations. In Humanoid tasks, ORCA(NP) achieves higher overall performance because TemporalOT almost entirely fails on every task. The effect of pretraining is most apparent when the demonstrations are temporally aligned because this setting satisfies the core assumption of TemporalOT. In Fig. 5, ORCA achieves an average normalized return of $0.57$ compared to ORCA(NP) ($0.33$) on aligned demonstrations.
@@
5.6 Does ORCA scale with more demonstrations?
Following the strategy of Fu et al. (2024c), ORCA can be adapted to multiple demonstration videos $\tilde{\Xi}=\{\tilde{\xi}^{1},\dots,\tilde{\xi}^{N}\}$ by calculating the ORCA reward at each timestep with respect to every video and max-pooling:
$\mathcal{R}_{\text{ORCA}}(o_{t},\xi,\tilde{\Xi})=\max_{\tilde{\xi}\in\tilde{\Xi}}\mathcal{R}_{\text{ORCA}}(o_{t},\xi,\tilde{\xi})$ (7)
To study how the number of demonstrations affects performance, we first investigate the setting where they all have the same speed, and the demonstrations only vary due to randomly initialized object locations. Figure 6 shows how ORCA’s performance improves as the number of demonstrations increases. In contrast, although TemporalOT handles multiple videos using the same method, and it initially benefits from having more than one demonstration, its performance starts to degrade given a large number of demonstrations.
We additionally train policies using 4 demonstrations with different speeds by randomly sampling one video from each of the four temporal misalignment categories described in Figure 4. ORCA outperforms TemporalOT in this scenario, as shown in figure 6. ORCA policies additionally perform better when trained on 4 demonstrations with different speeds than just a single demonstration ($0.67\pm 0.08\rightarrow 0.77\pm 0.08$), although they are still worse than 4 same speed demonstrations. Meanwhile, TemporalOT shows minimal gains with different speed demonstrations ($0.14\pm 0.02\rightarrow 0.16\pm 0.02$), demonstrating its inability to handle temporal misalignment at scale. Overall, ORCA consistently benefits from training with multiple demonstrations and exhibits greater robustness compared to baselines when these demonstrations have different speeds.
Figure 6: Results given more demonstrations. Same indicates that all demonstrations in the set are the same speed, whereas Diff indicates that they are different speeds. We report mean expert-normalized returns with standard error across 3 Meta-world tasks (Door open, Window open, Lever pull).
5.7 How does ORCA perform given a policy conditioned on image observations?
ORCA’s goal is to estimate rewards from a single demonstration in the observation space before learning a policy, making no assumptions about the policy itself. We additionally evaluate ORCA when the policy is conditioned on image observations instead of ground-truth states, following the DrQv2 setup of (Fu et al., 2024c). The results in Appendix C.5 demonstrate that the state-based and image-based policies trained with ORCA have similar performance (average normalized return: $0.704\pm 0.10\rightarrow 0.76\pm 0.06$). In contrast, policies trained with TemporalOT perform poorly, regardless of their input.
5.8 Visual Encoder Ablation
ORCA works with any visual encoder that can act as a distance metric between frames. We include in Appendix C.4 an additional ablation of ORCA with LIV (Ma et al., 2023), a robotics-specific visual encoder, and DINOv2 (Oquab et al., 2023), a standard vision model. The Resnet50 used for the main result in this paper achieves the best performance. However, there is high variability, substantiating the findings of prior work that evaluates different visual encoders for RL training (Hu et al., 2023). Ultimately, the best visual encoder is a practical and task-dependent choice.